OpenStudy (anonymous):

\[\int\frac{1}{\sqrt{1+x^{2}}} dx\] Without integration by trig. substitution

OpenStudy (anonymous):

how about using integral formula?

OpenStudy (apoorvk):

Well... Let \(x = \tan t\) then, \(dx = \sec^2tdt\) So, your integral becomes: \(\int \sec t dt\) which is back to @Callisto explanation in a post a few questions back. WHEEEEEEEEEEE!!!!!!!!!

OpenStudy (callisto):

Apoor.... NO TRIGO!!!!

OpenStudy (apoorvk):

I am BLIND. >.<

OpenStudy (anonymous):

PUT \[x=\tan \theta\]

OpenStudy (anonymous):

\[1+\tan ^2(\theta)=\sec^2(\theta)\]

OpenStudy (callisto):

@nitz According to the question: Without integration by trig. substitution

OpenStudy (anonymous):

sorry i didnt see it

OpenStudy (apoorvk):

The standard integrand formula is: \(\large \int \frac 1 {\sqrt{a^2 +x^2}} = ln|x+\sqrt{x^2 + a^2}| + k\) Now how do we derive this hmm...

OpenStudy (anonymous):

Without trig substitution. It's just a simple substitution and manipulation (:

OpenStudy (anonymous):

isnt it \[\tan^{-1} (x)\]

OpenStudy (anonymous):

no trig

OpenStudy (apoorvk):

You sure you do this WITHOUT that trig. thing?? Hmm..

OpenStudy (anonymous):

It's workable

OpenStudy (apoorvk):

@nitz that's for integral of 1/(1+x^2)

OpenStudy (anonymous):

i remember....its sumthing to do with hyperbolic functions.....

OpenStudy (anonymous):

I suppose we can't let\[x=\sinh(u)?\]

OpenStudy (apoorvk):

No @Mimi_x3 that sub. didn't work, back to square one.

OpenStudy (anonymous):

I don't know if this violates the trig substitution constraint but I'll post it anyway, maybe inspire some other thought. Let \[x=\sinh(u)\]\[dx=\cosh(u)\] So,\[\int\limits_{}\frac{1}{\sqrt{1+x^2}}=\int\limits_{}\frac{1}{\sqrt{1+sinh^2(u)}}cosh(u)=\int\limits_{}\frac{1}{\cosh(u)}\cosh(u)du=\int_{}du=u\] Must put back in terms of x. \[x=\sinh(u)\rightarrow u=\sinh^{-1}(x)\] So, \[\int\limits_{}\frac{1}{\sqrt{1+x^2}}dx=\sinh^{-1}(x)+c\]

OpenStudy (anonymous):

Supposed to be a du after the second integral...